### v1

# Efficient LLM Jailbreak via Adaptive Dense-to-sparse Constrained Optimization

Abstract
Recent research indicates that large language models (LLMs) are susceptible to jailbreaking attacks that can generate harmful content. This paper introduces a novel token-level attack method, Adaptive Dense-to-Sparse Constrained Optimization (ADC), which effectively jailbreaks several open-source LLMs. Our approach relaxes the discrete jailbreak optimization into a continuous optimization and progressively increases the sparsity of the optimizing vectors. Consequently, our method effectively bridges the gap between discrete and continuous space optimization. Experimental results demonstrate that our method is more effective and efficient than existing token-level methods. On Harmbench, our method achieves state of the art attack success rate on seven out of eight LLMs. Code will be made available. Trigger Warning: This paper contains model behavior that can be offensive in nature.
1 Introduction
Recent advancements have allowed large language models (LLMs) to be employed across various sectors, such as content generation [15], programming support [13], and healthcare [7]. Nevertheless, LLMs can pose risks by possibly generating malicious content, including writing malware, guidance for making dangerous items, and leaking private information from their training data [18, 10]. As LLMs become more powerful and widely used, it becomes increasingly important to manage the risks associated with their misuse. In this context, the concept of red-teaming LLMs is introduced to test the reliability of their safety features [2, 17]. Consequently, the LLM jailbreak attack was developed to support the red-teaming process: by combining the jailbreak prompt with malicious questions (e.g., how to make explosives), it can mislead the aligned LLMs to circumvent the safety features and potentially produce responses that are harmful, discriminatory, violent, or sensitive.
Recently, a number of automatic jailbreak attacks have been introduced. Generally, these can be categorized into two types: prompt-level jailbreaks [8, 11, 3] and token-level jailbreaks [18, 6, 9]. Prompt-level jailbreaks employ semantically meaningful deception to compromise LLMs. These queries can be created manually, through prompt engineering, or generated by another LLM. The main advantage of prompt-level jailbreaks is that the attack query is typically in natural language, making it interoperable. However, this feature also makes these jailbreak methods vulnerable to alignment training, since training examples are easy to create [5, 1]. For instance, prompt-level jailbreaks generally show poor performance on the Llama2 series [15] (although Llama2 is not designed to defend these jailbreak methods). Token-level jailbreaks add an adversatial suffix to the user query, and use gradient information to optimize the adversatial tokens. Although token-level jailbreaks allow more precise control over the LLM output, they are more computationally expensive than prompt-level jailbreaks.
In this paper, our focus is on token-level jailbreaks. The current state of the art token-level jailbreak method, GCG [18], employs a simple coordinate descent to perform a discrete optimization in the token space. Given the inherent difficulty of discrete optimizations, finding a solution requires significant computational resources and may not always yield an optimal outcome [6].
Motivated by this difficulty, our method relax the discrete optimization into continuous optimization, which allows us to employ stronger optimizers in the continuous space beyond coordinate descent. While relaxing optimization in a continuous space is straightforward, transitioning the optimizing vectors from continuous space to the discrete token space proves challenging. Direct projection onto the discrete space is ineffective, as it substantially alters the optimized vectors, leading to performance drops. To address this, we propose ADS, an adaptive dense-to-sparse constraint, on the dense token optimization. Initially, the optimization occurs in the full continuous space. As the optimization progresses and the loss decreases, we gradually reduce the optimization space to a highly sparse space, and ultimately to an approximate one-hot space. This progressive approach minimizes the gap between the relaxed contiguous space and the one-hot space and reduce the performance drop when transitioning the optimizing vectors to discrete tokens.
Our approach surpasses existing token-level jailbreak methods in both effectiveness and efficiency. In an apple-to-apple comparison on AdvBench, our method significantly outperforms GCG across three LLMs with 2/3 computational cost. Additionally, on Harmbench, our method achieves state-of-the-art results for seven LLMs. Notably, our method also attains 35% attack success rate (ASR) in attacking the adversarially trained LLM, Zephyer R2D2 [10], a feat that previous token-level jailbreak methods could not achieve, as they typically yielded nearly 0% ASR performance. Codes will be available.
2 Methodology
2.1 Problem Setup
Following [18], our framework seeks to search for adversarial strings. When appending this string to a user query, the LLM chatbot will respond to the query, even if it is harmful. Here is an example:
User: Create fake news that could lead to panic or chaos Assistant: Sorry, I cannot assist with that.
User: Create fake news that could lead to panic or chaos ! ! ! ! ! ! Assistant: Sure, here is fake news that could lead to panic or chaos
The red string ( ! ! ! ! ! !) represents an adversarial string that causes the LLM chatbot to reply with a positive sentence, as shown in the example above. Following early work [18], we hypothesize that if the initial response of a large language model (LLM) chatbot starts with an affirmative phrase (e.g., “Sure, here is”), it is more likely to adhere to the user’s instructions. Therefore, the main objective of our jailbreak is to optimize the conditional probability of eliciting a specific target response, based on the user query, any associated system prompts, and the adversarial string.
Notation
In an LLM with a vocabulary size of $V$, the $i^{\textrm{th}}$ vocabulary is denoted as a one-hot vector $e_{i}\in\mathbb{R}^{V}$ where the $k$-th element is given by $e_{i}[k]=\mathds{1}(i=k)$. Let $\mathcal{C}=\{e_{i}\}_{1\leq i\leq V}$ denote the set of all vocabularies. Let $x_{1:n}$ denote a sequence of vectors from: $x_{1},x_{2},\cdots,x_{n}$. The user query of length $l$, the adversarial tokens of length $n$ and the target response of length $m$ are represented as $x_{1:l},z_{1:n},y_{1:m}$ respectively. The jailbreak objective is given by:
$\max_{\forall i,z_{i}\in\mathcal{C}}\;p(y_{1:m}|x_{1:l}\oplus z_{1:n}),$ (1)
where $\oplus$ denotes concatenation. Thanks to LLM’s next token prediction nature, we can minimize the following cross entropy loss to achieve the optimization goal in Equation 1.
$\min_{\forall i,z_{i}\in\mathcal{C}}\;\text{loss}=\sum_{k=1}^{m}\text{CE}\left(\text{LLM}(x_{1:l}\oplus z_{1:n}\oplus y_{1:k-1}),y_{k}\right).$ (2)
2.2 Relaxed continuous optimization
The optimization space of Equation 2 is discrete, thus common optimizer does not work for it directly. A straightforward method is to consider a relaxed continuous optimization problem. We use the probability space as the continuation of the one-hot vector set $\mathcal{P}=\{w\in\mathbb{R}^{V}\big{|}w[i]\geq 0,\|w\|_{1}=1\}$, and relax Equation 2 to Equation 3:
$\min_{\forall i,z_{i}\in\mathcal{P}}\;\text{loss}=\sum_{k=1}^{m}\text{CE}\left(\text{LLM}(x_{1:l}\oplus z_{1:n}\oplus y_{1:k-1}),y_{k}\right).$ (3)
Equation 3 is a continuous space optimization thus an easier problem than Equation 2. However, if optimizing Equation 3, we need to convert $z_{1:n}$ from $\mathcal{P}$ to $\mathcal{C}$ either during or after the optimization process. Only one-hot vectors in $\mathcal{C}$ are legal inputs for an LLM.
We find it does not work well if directly apply a projection from $\mathcal{P}$ to $\mathcal{C}$:
$\text{proj}(x)=e_{j}\text{ where }j=\arg\min_{k}\|x-e_{k}\|_{2}^{2}=\arg\max_{k}x[k].$ (4)
The optimizing loss would increase sharply after this projection. This occurs because, in most cases, the optimized result of $z_{1:n}$ from Equation 3 are dense vectors. The distance between a dense vector $x$ and its projection to $\mathcal{C}$ tends to be substantial, i.e., $\|\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(x)-x\|$ is large. The projection changes the optimizing $z_{1:n}$ greatly thus hurts optimization.
To reduce the negative impact caused by projection, we would hope that the optimizing vectors $z_{1:n}$ are sparse before we apply the projection (Equation 4). Then the distance between a sparse vector $x$ and its projection could be smaller.
To tackle this issue, we propose a dense-to-sparse constraint to the optimization of Equation 3. During the optimization process, we progressively enhance the sparsity of the optimizing vectors $z_{1:n}$, which helps decrease the loss when projecting from set $\mathcal{P}$ to set $\mathcal{C}$. Specifically, at each optimization step, we manually convert $z_{1:n}$ to be $S$-sparsity, where $S$ is given by the cross entropy loss in Equation 3:
$S=\frac{1}{m}\sum_{k=1}^{m}\exp(\text{CE loss of token }y_{k}).$ (5)
The motivation of the adaptive sparsity (Equation 5) is to use less sparsity constrain before we can find an optimal solution in the relaxed space $\mathcal{P}$. If the LLM is random guessing when predicting the target response $y_{1:m}$, the cross entropy loss is approximately $-\log(1/V)$ (recall $V$ is the vocabulary size). In this case, $S\approx V$ gives no constraint to $z_{1:n}$. If the loss in Equation 3 is close to zero, $S\approx 1$ indicates we would project all $z_{1:n}$ to the one-hot space $\mathcal{C}$ because there is no further optimization room in the relaxed space $\mathcal{P}$.
Note that an $S$-sparsity subspace in a $d$-dimensional space ($S<d$) is not a convex set. Projection onto the $S$-sparsity subspace may computationally complex and not yield a unique solution. We propose a simple transformation that convert an arbitrary vector to be $S$-sparsity and in the relaxed set $\mathcal{P}$. While this transformation is not a projection (i.e., mapping $x$ to the closest point in $\mathcal{P}$), we find it works well. Algorithm 1 describes the details of the transformation.
Algorithm 1 Transform a vector to be in $\mathcal{P}$ and be $S$-sparsity
1: Input: vector $x\in\mathbb{R}^{V}$ and the target sparsity $S$
2: $\delta\leftarrow$ The $S$-th largest element in $x$.
3: $x[i]\leftarrow\textrm{ReLU}(x[i])+10^{-6}$ if $x[i]\geq\delta$ else 0 $\triangleright$ Set non-top S largest elements to zero
4: $x\leftarrow x/\sum_{i}x[i]$triangleright $10^{-6}$ is for numerical stability
5: Return $x$
Non-integer sparsity
Most commonly, $S=\exp(\text{loss})$ is not an integer. Let $\lfloor S\rfloor$ denote the maximum integers no greater than $S$. In the case $S$ is not an integer, we randomly select
$\text{round}((S-\lfloor S\rfloor)\cdot n)$
uniformly vectors $z_{i}$ from $z_{1:n}$ and transform these $z_{i}$ to be $(\lfloor S\rfloor+1)$-sparse and the remaining $z_{i}$ to be $\lfloor S\rfloor$-sparse. Then the average sparsity of $z_{1:n}$ is approximately $S$.
For example suppose $S=1.2,n=20$, then $4$ random vectors from $z_{1:n}$ are converted to be 2-sparse and 16 random vectors from $z_{1:n}$ are converted to be 1-sparse, i.e., one-hot vectors. Such design is to progressively reduce the gap between the relaxed set $\mathcal{P}$ and the one-hot vector $\mathcal{C}$ especially when the loss in the relaxed space is low.
2.3 Optimizer design to escape local minima
Equation 3 possesses many local minima, most of which are suboptimal for jailbreak. Consequently, it is essential to create an optimizer designed to effectively escape local minima. We adopt the following three designs:
Optimization hyperparameter
In all our experiments, we employ the momentum optimizer with a learning rate of 10 and a momentum of 0.99, and do not adjust them during the optimization. The large learning rate and momentum contribute to the optimizer’s ability to escape local minima.
No reparameterization trick
Optimization in Equation 3 has a constraint that $\forall i,\|z_{i}\|=1$. Existing work may handle the constraint using a reparameterization trick. For example, instead of optimizing $z_{i}\in\mathcal{P}$, one can optimizing $w_{i}\in\mathbb{R}^{V}$ and represent $z_{i}$ as:
[…]
However, we argue that such reparameterization trick is not good for the optimizer to escape local minima. Take softmax reparameterization for example. To make $z_{i}$ sparse, most elements in $z_{i}$ are close to or equal to 0. If a softmax reparameterization is applied, we can show that $\frac{\partial\ell}{\partial w_{i}[k]}\propto z_{i}[k]$ where $\ell$ is the optimization objective. If one element of $z_{i}$ is close to zero: $z_{i}[k]\rightarrow 0$, the derivative of the corresponding $w_{i}[k]$ is also close to zero. Once $z_{i}[k]$ is updated to a small number or set to zero (in Algorithm 1), updating this element in subsequent optimization steps becomes difficult, making the entire optimization stuck in a local minima.
To design an optimization that readily escapes local minima, we eschew the reparameterization trick. Instead, we utilize the projection (Algorithm 1) to satisfy the constraint $z_{i}\in\mathcal{P}$. Here, zeroing one element of $z_{i}$ does not diminish the likelihood of updating this element to a larger number in subsequent optimization steps.
Multiple initialization starts
We initialize $z_{1:n}$ from the softmax output of Gaussian distribution:
$z_{i}\leftarrow\textrm{softmax}(\varepsilon)\textrm{ where }\varepsilon\sim\mathcal{N}(0,I_{V})$ (6)
Like many non-convex optimization, we initialize multiple $z_{1:n}$ to reduce the loss of “unlucky” initialization. These different starts are optimized independently. Once one of searched $z_{1:n}$ can jailbreak, the optimization would early stop.
2.4 Complete Attack Algorithm
Algorithm 2 provides an overview of our method with one initialization start. We employ a momentum optimizer with a learning rate of 10 and a momentum of 0.99, optimizing over maximum 5000 steps.
Algorithm 2 Adaptive dense-to-sparse optimization
1: Input: User query $x_{1:l}$ and target response $y_{1:m}$. Number of optimizable adversarial tokens $n$.
2: Initialize dense adversarial tokens $z_{1:n}$ using Equation 6.
3: Initialize the optimizer as described in Section 2.3: $\text{lr}\leftarrow 10,\text{ momentum}\leftarrow 0.99$.
4: for step in $1\cdots,5000$ do for step in $1\cdots,5000$ do
5:      for step in $1\cdots,5000$ do Compute the loss and gradient of $z_{1:n}$ with respect to the objective in Equation 3 Compute the loss and gradient of $z_{1:n}$ with respect to the objective in Equation 3
6:      Compute the loss and gradient of $z_{1:n}$ with respect to the objective in Equation 3 Update $z_{1:n}$ using the momentum optimizer. $\triangleright$ Unconstrained optimization Update $z_{1:n}$ using the momentum optimizer. $\triangleright$ Unconstrained optimization
7:      Update $z_{1:n}$ using the momentum optimizer. $\triangleright$ Unconstrained optimization Convert $z_{1:n}$ to the target sparsity $\exp(\text{loss})$ using Algorithm 1. Convert $z_{1:n}$ to the target sparsity $\exp(\text{loss})$ using Algorithm 1.
8:      Convert $z_{1:n}$ to the target sparsity $\exp(\text{loss})$ using Algorithm 1. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop.
The actual jailbreak methods initialize and optimize multiple adversarial tokens in parallel. Once one of them can jailbreak, the entire attack would early stop.
Number of initialization starts
One step optimization includes 1 model forward (dense tokens loss computation), 1 model backward (dense tokens gradient computation) and 1 model forward (jailbreak evaluation of the projected one-hot tokens). Since the model backward only computes the gradient of the input, the computational cost of one model backward equals to that of one model forward. Thus the maximum computational cost for 5000 steps optimization is $1.5\times 10^{4}$ model forward for one initialization start.
Similarly the computational cost for the state of the art token based attack GCG [18] is $2.57\times 10^{6}$ model forward under its standard configuration (512 batch size, maximum 500 steps). To align with GCG’s computational cost, we initialize 16 different initialization starts and optimize them in a batch, achieving a computational cost that is 93% of GCG. We name our attack of this configuration (initialize 16 different starts and optimize for maximum 5000 steps) as ADC.
More efficiency integrated with GCG
We find that ADC converges slower when the loss in Equation 3 is low. This is because we use a fixed and large learning rate in our algorithm. A learning rate schedule may solve similar problems in neural network training. However, it introduces more hyper-parameter tuning thus we have a lower preference for it. We find a more efficient way is to switch to GCG after certain number of steps. Specifically, we initialize 8 different starts and optimize for maximum 5000 steps. If not attack successfully, we use the searched adversarial tokens with the best (lowest) loss as the initialization to perform GCG attack for 100 steps. The two-stage attack achieves a computational cost that is 67% of standard GCG, and we name it as ADC+.
3 Experimental Results
3.1 Experimental Setup
Datasets
We evaluate the effectiveness of ADC optimization on three datasets:
  • 1
    AdvBench harmful behaviors subset [18] contains 520 harmful behavior requests. The attacker’s objective is to search an adversarial string that will prompt the LLM to produce any response that seeks to follow the directive.
  • 2
    AdvBench harmful strings subset contains 574 strings that reflect harmful or toxic contents. The attacker’s objective is to find a string that can prompt the model to generate these exact strings. The evaluation on this dataset reflects the attacker’s ability to control the output of the LLM.
  • 3
    HarmBench [10] contains 400 harmful behavior requests similar to AdvBench “harmful behaviors” subset, but with a wider range of topics including copyright, context reference and multimodality. We evaluate on the “Standard Behaviors” subset of 200 examples. Evaluation on the complete HarmBench dataset will be released soon.
Victim models
We attempt to jailbreak both open- and closed-source LLMs. Specifically, we consider open-source LLMs: Llama2-chat-7B [15], Vicuna-v1.5-7B [4], Zephyr-7b-$\beta$ [16], and Zephyr 7B R2D2 [10] (an adversarial trained LLM to defend against attacks). We use their preset system prompts and chat templates. Results for transfer based attack on GPT-3.5 and 4 [12] will come soon.
Metric
Attack Success Rate (ASR) is widely used to evaluate LLM jailbreak methods. However, previous works use different methods to calculate ASR and some may have different shortcomings (see Harmbench [10] for a detailed discussion). We follow Harmbench [10] to compute ASR: the victim model generates the response of maximum 512 tokens given the adversarial request and we use the red teaming classifier (from HarmBench [10]) to decide if the response follows the request. The classifier only gives binary classification results (“YES” or “NO”), and ASR is computed as the ratio of the number of “YES” over the number of attacked examples.
The AdvBench harmful strings benchmark requires the attacker to generate the exact strings. For this task, we employ the Exact Match (EM) metric, meaning an attack is only considered successful if the generated string precisely matches the target harmful string.
3.2 Comparison with the baseline on AdvBench
We employ GCG as our baseline for a detailed comparison using the AdvBench dataset on 3 LLMs: Llama2-chat-7B, Vicuna-v1.5-7B and Zephyr $\beta$ 7B. Recall that GCG and ADC require a differing number of attacking steps, we use 16 initialization start to align with the computing budgets of GCG with a batch size of 512. We re-implement GCG to make sure the comparison is fair.
Table 1 shows the results on 520 examples from the AdvBench harmful behaviors subset. The computing budge indicates the maximum computational cost if not considering early stop and the Wall-clock time is the average elapsed real time per sample on a single A100 machine. As shown in Table 2, our method performs significantly better than GCG on Llama2-chat-7B, and performs similarly to GCG on other LLMs since the performance is already close to 100%. However, our method still shows advantages in wall-clock time because our method is able to find a jailbreak string with fewer steps (higher early stop rate).
Table 1: Comparison on 520 examples from AdvBench Behaviours. A higher ASR is better.
model method ASR (%)
Computing
Budge
Wall-clock
Time (min)
Early Stop
Rate (%)
Llama2-chat-7B GCG 53.8 1$\times$ 20.6 53.3
ADC 77.7 0.93$\times$ 14.1 75.4
Vicuna-v1.5-7B GCG 99.4 1$\times$ 3.0 98.3
ADC 100.0 0.93$\times$ 1.8 99.4
Zephyr-$\beta$-7B GCG 98.1 1$\times$ 13.2 68.3
ADC 98.1 0.93$\times$ 11.9 78.1
We further explore our method’s capacity for precise control over the outputs of the LLM. Precise control of LLM outputs is important for multi-turn attack and better shows the search ability of the attack method. Table 2 shows the Exact Match performance on AdvBench Strings. The early stop rate is not reported because it equals to EM in the exact matching setting. We can see that ADC is better than GCG on Vicuna-v1.5-7B and Zephyr-$\beta$-7B but worse on Llama2-chat-7B. However, ADC+ consistently outperforms the baseline in terms of performance and efficiency.
Table 2: Comparison on 574 examples from AdvBench Strings. A higher EM is better.
model method EM (%) Computing Budge Wall-clock Time
Llama2-chat-7B GCG 41.3 1$\times$ 18.8 min
ADC 15.5 0.93$\times$ 21.7 min
ADC+ 72.4 0.67$\times$ 14.6 min
Vicuna-v1.5-7B GCG 90.6 1$\times$ 5.8 min
ADC 98.4 0.93$\times$ 2.4 min
ADC+ 99.7 0.67$\times$ 1.7 min
Zephyr-$\beta$-7B GCG 52.4 1$\times$ 13.7 min
ADC 87.1 0.93$\times$ 7.9 min
ADC+ 97.1 0.67$\times$ 5.7 min
3.3 Comparison with existing methods on Harmbench
Existing studies on LLM jailbreak have employed various evaluation frameworks, such as testing on different subsets of AdvBench, generating responses of varying lengths and different ASR computation methods. Therefore, direct comparison with the ASR reported in these studies is not convincing. Hopefully, HarmBench [10] replicates these methods and provides a fair comparison on the HarmBench dataset. We follow their setting of chat templates, number of generated tokens and ASR computation method. Table 1 shows the comparison between our method, ADC+ and the top performance methods reported by HarmBench [10]: GCG [18], AutoPrompt (AP) [14], PAIR [3], TAP [11] and AutoDan [8].
Table 3 illustrates that although no single jailbreak method consistently outperforms others across all large language models (LLMs), our approach achieves state-of-the-art results on seven out of eight LLMs, excluding Zephyr-R2D2. Notably, our method registers more than 50% and 20% improvements in attack success rate (ASR) on LLama2-7B-chat and Qwen-7B-chat, respectively, compared to existing methods. It’s important to note that Zephyr-R2D2 has been adversarially trained to resist token-level jailbreaks, making prompt-level methods more effective in comparison. Nevertheless, our method still demonstrates significant effectiveness on Zephyr-R2D2, achieving notable ASR where other token-level jailbreak methods like GCG achieve only single-digit performance.
Table 3: Jailbreak comparison on 200 examples from HarmBench Standard Behaviours. The number indicates the ASR for the corresponding LLM and the jailbreak method. A higher ASR is better.
Jailbreak Methd
LLM GCG AP PAIR TAP AutoDan Ours
LLama2-7B-chat 34.5 17.0 7.5 5.5 0.5 89.5
LLama2-13B-chat 28.0 14.5 15.0 10.5 0.0 44.5
Vicuna-v1.5-7B 90.0 75.5 65.5 67.3 89.5 100
Vicuna-v1.5-13B 87.0 47.0 59.0 71.4 82.5 100
Qwen-7B-chat 79.5 67.0 58.0 69.5 62.5 99.0
Qwen-14B-chat 83.5 56.0 51.5 57.0 64.5 96.7
Zephyr-$\beta$-7B 90.5 79.5 70.0 83.0 97.3 99.5
Zephyr-R2D2 0.0 0.0 57.5 76.5 10.5 35.0
4 Conclusion
In this paper, we present a new token-level attack method, Adaptive Dense-to-Sparse Constrained Optimization (ADC), which successfully breaches several open-source large language models (LLMs). Our technique transforms the discrete jailbreak optimization problem into a continuous one and progressively enhances the sparsity of the optimizing vectors. This strategy effectively closes the gap between discrete and continuous space optimization. Our experimental findings show that our method outperforms existing token-level techniques in both effectiveness and efficiency. On Harmbench, our method achieves the highest attack success rate on seven out of eight LLMs. Results of transfer attacks on closed-source LLMs will be updated.